# HiMaCon: Discovering Hierarchical Manipulation Concepts from Unlabeled Multi-Modal Data

### v1

@@ Appendix B Pseudocode
Appendix C More Study on Learned Manipulation Concepts
C.1 Additional Experiments on Enhanced Imitation Learning
Data Constraint Experiments
L90-90 50 demos/task 25 demos/task 10 demos/task
Ours 89.6 ± 0.6 77.6 ± 0.5 61.2 ± 1.1
Plain 75.1 ± 0.6 70.1 ± 0.3 59.1 ± 0.9
Future Prediction Strategy
L90-90 Ours Next-n Next-random
DP 89.6 ± 0.6 83.0±0.3 82.8 ± 0.4
C.2 Alignment with Semantic Sub-Goals
Fig. 8 shows results from analyzing demonstrations from three tasks:
  • Task #1: Open the top drawer of the cabinet and put the bowl in it;
  • Task #2: Close the bottom drawer of the cabinet and open the top drawer;
  • 1.
  • 2.
  • 3.
In each heatmap, the entry at position $(i,j)$ represents the average similarity between categories $i$ and $j$. For readability, only the top three similarity values in each row are displayed.
C.3 Motion Study
We evaluate whether the learned manipulation concept latents capture the robot’s motion. Using Eq. 10, we calculate the average similarity between movements based on manipulation concept latents corresponding to specific gripper actions. Specifically, we collect latents for the following movements from task demonstrations in LIBERO-90:
  • 1.
  • 2.
  • 3.
  • 4.
C.4 Diversity & Discrimination Study
C.5 Multi-Level Hierarchical Structure
C.6 Real Robot Experiments Details
C.7 Multi-Horizon Goal Prediction Visualization
Dataset.
Manipulation Concept Discovery.
Appendix D Limitations & Future works
Table 9: Comparison of Diffusion Policy enhanced with our manipulation concepts versus state-of-the-art vision-language-action models.
LIBERO-LONG LIBERO-GOAL
$\pi_{0}$ + FAST (fine-tuned) [41] 60.2 88.6
$\pi_{0}$ (fine-tuned) [4] 85.2 95.8
OpenVLA-OFT [22] 94.5 97.9
DP (Enhanced by Manipulation Concepts) 89.0 95.7
Scaling up. Computational constraints have limited our exploration of how our method scales with larger datasets. We plan to leverage pretrained multi-modal foundation models, adopting structures inspired by [7] and extending pretraining beyond robotics data as in [62]. We also aim to investigate whether our manipulation concepts can enhance advanced policies like Vision-Language-Action models [4, 18, 22, 23]. Tab. 9 suggests that our concept-enhanced policies achieve comparable performance to these advanced models despite requiring substantially less training data, indicating significant potential for improving learning efficiency.

### v2

@@ Appendix B Pseudocode
Appendix C More Study on Learned Manipulation Concepts
C.1 Additional Experiments on Enhanced Imitation Learning
Sampling Strategies
In this part, we focus on methodology for deriving hierarchical structures from learned representations (Sec. 3.3). While we adopt a threshold-based hierarchy derivation method (Eq. 4) as a proof of concept, we acknowledge that alternative derivation methodologies warrant further investigation (see Sec. D). For the threshold-based approach, we employ uniform sampling of the threshold $\epsilon$ during training. This choice ensures full coverage of all possible hierarchical structures, as we do not know a priori which threshold values might be suboptimal. To validate this design choice, we conduct an ablation study comparing different sampling strategies for $\epsilon$ in Eq. 7:
Table 7: Sampling Strategies Ablation. We compare different sampling strategies for $\epsilon$ in Eq. 7. Manipulation concepts are learned from LIBERO-90 and applied to policy learning on LIBERO-90. We report success rates (%).
Sampling Strategy Description ACT DP
Uniform (Ours) $\epsilon\sim\mathcal{U}\big(0,1\big)$ 74.8±0.8 89.6±0.6
Sparse $\epsilon\sim\{0.1,0.2,\cdots,1.0\}$ 67.6±0.5 81.1±0.8
Biased $\epsilon\sim\mathcal{U}\big(\frac{1}{3},\frac{2}{3}\big)$ 65.6±0.7 78.7±0.4
As shown in Tab. 7, uniform sampling currently achieves the best performance across both policy architectures. We hypothesize that while task-specific sampling strategies might excel on particular subsets, uniform sampling provides robust performance across diverse tasks due to its comprehensive coverage of the threshold space. Future work could explore adaptive sampling strategies tailored to specific task distributions.
Learning Methodology Contribution
We conduct an ablation study to isolate the contributions of our two core learning methodologies: Capturing Multi-Modal Correlations (Sec. 3.2) and Representing Multi-Horizon Sub-Goals (Sec. 3.3). Tab. 8 compares three configurations: (1) Cross-modal only: learning with only cross-modal alignment objectives in Eq. 3, (2) Multi-horizon only: learning with multi-horizon sub-goal prediction in Eq. 7 but without cross-modal alignment, and (3) Full method: combining both cross-modal alignment and multi-horizon prediction.
Table 8: Methodology Contribution Ablation. We evaluate the contribution of each learning component by training manipulation concepts on LIBERO-90 and applying them to policy learning on LIBERO-90. We report success rates (%).
Method ACT DP
Cross-modal only 69.1±0.6 82.8±1.0
Multi-horizon only 71.6±0.4 80.5±0.5
Ours (Full method) 74.8±0.8 89.6±0.6
The results in Tab. 8 reveal that both components make substantial and complementary contributions to performance. We attribute this synergy to the distinct roles of each component: cross-modal alignment grounds the understanding of correlations across different modalities, while multi-horizon prediction captures hierarchical temporal structure. Together, they enable the learning of manipulation concepts that are both correlationally coherent and temporally structured, leading to more robust policy learning.
Data Constraint Experiments
50 demos/task 25 demos/task 10 demos/task
Ours 89.6 ± 0.6 77.6 ± 0.5 61.2 ± 1.1
Plain 75.1 ± 0.6 70.1 ± 0.3 59.1 ± 0.9
Distance Metric
We conduct an ablation study comparing spherical distance and cosine distance $\frac{1-\cos{(\cdot})}{2}$ for $\text{dist}(\cdot,\cdot)$ in Eq. 4. Tab. 10 reports the performance when concepts are learned and applied to LIBERO-90 tasks. Further investigation into distance-threshold-based subprocess derivation methods represents a promising direction for future work.
Table 10: Ablation study on distance metrics for concept learning on LIBERO-90. Spherical distance consistently outperforms cosine distance across both baseline methods.
Cosine Distance Spherical Distance (Ours)
ACT 67.8±0.5 74.8±0.8
DP 82.0±0.4 89.6±0.6
Sub-process Derivation
We conduct an ablation study comparing two approaches for constraining manipulation concept latents within each sub-process in Eq. 4. Our proposed method enforces proximity among all concept latents throughout the sub-process (“Sequential Constraint”), while the baseline only constrains the distance between the initial and final concept latents (“Endpoint Constraint”). We evaluate both approaches on LIBERO-90, where concept discovery and policy enhancement are performed. Tab. 11 reports the task success rates when integrating the learned manipulation concepts with different policy architectures.
Table 11: Ablation study on sub-process derivation constraints. We compare enforcing proximity among all manipulation concept latents within each sub-process (Sequential Constraint) versus constraining only the initial and final latents (Endpoint Constraint). Results show average success rates (%) with standard errors across LIBERO-90 tasks.
Sequential Constraint Endpoint Constraint
ACT 74.8±0.8 68.4±0.8
DP 89.6±0.6 79.8±0.5
Future Prediction Strategy
L90-90 Ours Next-n Next-random
DP 89.6 ± 0.6 83.0±0.3 82.8 ± 0.4
Usage of Manipulation Concept Encoder
We investigate two strategies for leveraging the manipulation concept encoder from Eq. 1 in downstream policy learning. The encoder serves as an intermediate module that extracts manipulation concept representations from demonstrations. We compare the following approaches: (1) Direct Conditioning: The trained encoder directly processes current observations to generate manipulation concepts, which are then concatenated with observations as additional input features to the policy network. (2) Joint Prediction (Ours): The policy network is trained to jointly predict both future actions and future manipulation concepts from current observations, as described in Sec. 3.4. Tab. 13 presents the comparative results across two policy architectures.
Table 13: Comparison of Manipulation Concept Usage Strategies.
Policy Direct Conditioning Joint Prediction (Ours)
ACT 71.1±0.4 74.8±0.8
DP 79.3±0.9 89.6±0.6
The performance gap stems from a temporal alignment mismatch between concept representations and action predictions. In Direct Conditioning, the encoder extracts concepts from current or past observations, creating a temporal lag: the policy receives historical concept information when planning future actions. In contrast, Joint Prediction enforces temporal coherence by training the policy to predict future manipulation concepts alongside future actions, ensuring that the predicted concepts align temporally with the planned action sequence.
This temporal alignment is critical in multi-phase manipulation tasks. For example, consider a pick-and-place scenario: immediately after grasping an object, the current observation encodes grasping-related dynamics. However, to execute the subsequent placement action, the policy requires placement-relevant information. Joint Prediction learns to anticipate these future task-phase concepts, providing the policy with forward-looking contextual information. Direct Conditioning, by contrast, conditions the policy on backward-looking grasping concepts that offer limited guidance for placement planning.
While our results demonstrate the advantages of temporal alignment through joint prediction, we acknowledge that direct conditioning on historical concepts may benefit tasks requiring explicit long-horizon memory or reactive behaviors based on past states [14]. Future work will explore hybrid architectures that combine both strategies.
C.2 Alignment with Semantic Sub-Goals
Fig. 9 shows results from analyzing demonstrations from three tasks:
  • Task #1: Open the top drawer of the cabinet and put the bowl in it;
  • Task #2: Close the bottom drawer of the cabinet and open the top drawer;
  • 1.
  • 2.
  • 3.
In each heatmap, the entry at position $(i,j)$ represents the average similarity ($\times 10.0$) between categories $i$ and $j$. For readability, only the top three similarity values in each row are displayed.
C.3 Motion Study
We evaluate whether the learned manipulation concept latents capture the robot’s motion. Using Eq. 10, we calculate the average similarity ($\times 100.0$) between movements based on manipulation concept latents corresponding to specific gripper actions. Specifically, we collect latents for the following movements from task demonstrations in LIBERO-90:
  • 1.
  • 2.
  • 3.
  • 4.
C.4 Diversity & Discrimination Study
C.5 Multi-Level Hierarchical Structure
C.6 Real Robot Experiments Details
C.7 Multi-Horizon Goal Prediction Visualization
Dataset.
Manipulation Concept Discovery.
C.8 Preliminary VLA Integration
We present a preliminary exploration of integrating manipulation concepts with vision-language-action models (VLAs). We build upon OpenVLA-OFT [23], which fine-tunes OpenVLA using pretrained parameters and a novel action adapter for downstream tasks. The action adapter processes hidden layer features from the original pretrained VLA model. Following this architecture, we introduce an additional “concept adapter” that implements the method described in Sec. 3.4, enabling the integration of manipulation concepts into the VLA.
To evaluate the data efficiency gains from manipulation concepts, we fine-tune the enhanced VLA on 50% of the training data used for LIBERO-10 tasks in the original OpenVLA-OFT study [23]. We compare fine-tuning performance with and without manipulation concept integration. Fig. 8 presents the results, where the x-axis indicates training epochs and the y-axis shows success rates for checkpoints at each epoch. The solid lines labeled “best” represent the highest success rate achieved up to that epoch.
The results demonstrate that manipulation concepts improve data utilization. With only half the training data, the concept-enhanced approach consistently achieves higher success rates throughout training. Notably, the original OpenVLA-OFT achieved 94.5% success with the full dataset [23], while our concept-enhanced model with half the data reaches comparable performance levels, indicating substantially improved data efficiency.
Figure 8: Data efficiency comparison on LIBERO-10 tasks with 50% training data. Solid lines show best performance up to each epoch for models with and without manipulation concepts.
We hypothesize that this improvement stems from HiMaCon’s ability to capture manipulation dynamics at multiple abstraction levels. The learned concepts provide explicit intermediate representations that bridge high-level task instructions and low-level control actions, thereby reducing the learning burden on VLAs by supplying structured manipulation knowledge rather than requiring learning of complex sensorimotor patterns from scratch. Further investigation of this integration will be pursued in future work.
Appendix D Limitations & Future works
Scaling up. Computational constraints have limited our exploration of how our method scales with larger datasets. We plan to leverage pretrained multi-modal foundation models, adopting structures inspired by [7] and extending pretraining beyond robotics data as in [68]. We also aim to further investigate whether our manipulation concepts can enhance advanced policies like Vision-Language-Action models [4, 19, 23, 24].